OpenStudy (anonymous):

sqrt24-3sqrt45+2sqrt20

OpenStudy (anonymous):

don't post same question over and over.

OpenStudy (mertsj):

\[\sqrt{24}-3\sqrt{45}+2\sqrt{20}\] \[\sqrt{4(6)}-3\sqrt{9(5)}+2\sqrt{4(5)}\] \[2\sqrt{6}-9\sqrt{5}+4\sqrt{5}\] \[2\sqrt{6}-5\sqrt{5}\]